OpenStudy (anonymous):

The value of 2*5^2 is 100.

OpenStudy (anonymous):

no it's 50

OpenStudy (anonymous):

You have to use PEMDAS Solve the exponents first 2(5)^2 = 2(25) = 50

OpenStudy (anonymous):

Squaring a number is the same as multiplying the number by itself

HanAkoSolo (jamierox4ev3r):

The people are correct… it is 50… the reason why is because…. in the Order of Operations, Square roots are first. 5^2 =25 and 25 multiplied by 2 equals 50. listen to the ambassador, will ya? xDD

HanAkoSolo (jamierox4ev3r):

if you still don't understand, let anyone of us know